OpenStudy (anonymous):

A rancher has 400 feet of fencing with which to enclose two adjactent rectangular corrals. What dimensions should be used so that enclosed area will be maximum?

OpenStudy (anonymous):



OpenStudy (anonymous):

I know that the Area of a rectangle is A=l*w

OpenStudy (anonymous):

I mean Area=2x*y

OpenStudy (anonymous):

if i had an area already i could solve for x or y and substitute back in the formula and solve for one then the other but I don't know how to start since i only have 400ft of fence to start with

OpenStudy (anonymous):

i know it can't be that hard but i am stuck on how to start out

OpenStudy (anonymous):

Area = y*2x, Fence Length: 4x + 3y = 400

OpenStudy (anonymous):

if the corals are adjacent, why put a fence between them? we get perimeter = \(4x+2y=400\)

OpenStudy (anonymous):

I am not a farmer, however, one rational is to be able to separate live stock.

OpenStudy (anonymous):

in the diagram in the book the diagram shows them seperated too

OpenStudy (anonymous):

x = 50 and y = 200/3 You should verify these answers yourself.